OpenStudy (anonymous):

Use the triangle at the right. Find the length of the missing side. 1. a = 16, b = 63 2. b = 2.1, c = 2.9

OpenStudy (anonymous):

I'm going to assume that these are the measures of the lengths. To find the missing lengths, you're going to use Pythagorean theorem. Pythagorean Theorem: \[a^2+b^2=c^2 \] So, for number 1, it'd look like this:\[16^2+63^2=c^2\]\[256+3969=\sqrt{4225}\]\[c^2=\sqrt{4225}\]\[c=65\] Using this, can you solve number 2? :o)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no